OpenStudy (anonymous):

f(x) = sqrt(x+9); g(x) = 8x - 13 Find f(g(x))

OpenStudy (anonymous):

Am I putting the g(x) in f(x) or vice versa? I know how to solve afterwards

OpenStudy (swissgirl):

Your g(x) becomes then new x value f(8x-13)

OpenStudy (anonymous):

f(g(x))= 8sqrtx-4?

OpenStudy (swissgirl):

\(f(x)= \sqrt{x+9}\) Now since its f(g(x)) That is another way of saying f(8x-13). Meaning we replace our x with 8x-13 \( f(8x-13)= \sqrt{(8x-13)+9 }\) Now we have to simplify by opening the brackets \( f(8x-13)= \sqrt{(8x-4) }\) \( f(8x-13)= \sqrt{4(2x-1) }\) \( f(8x-13)= 2\sqrt{2x-1 }\)

OpenStudy (anonymous):

oh ok I think I got it, Im going to try it on my last problem now. thanks so mucH!

OpenStudy (swissgirl):

No problem. You can always just come back here and check if you got your answer correct :)